# Improving Graph Neural Network Training Efficiency By Using Top Non-Robust Samples In The Training Set

### v3

@@ Abstract
Graph Neural Networks (GNNs) are a highly effective neural network architecture for processing graph-structured data. Unlike traditional neural networks that rely solely on the features of the data as input, GNNs leverage both the graph structure, which represents the relationships between data points, and the feature matrix of the data to optimize their feature representation. This unique capability enables GNNs to achieve superior performance across various tasks. However, it also makes GNNs more susceptible to noise from both the graph structure and the data features, which can significantly degrade their performance in common tasks such as classification and prediction. To address this issue, this paper proposes a novel method for constructing training sets by identifying training samples that are particularly sensitive to noise for a given model. These samples are then used to enhance the model’s ability to handle noise-prone instances effectively. Experimental results demonstrate that this approach can significantly improve training efficiency.
Index Terms:
Graph Neural Networks, Training, Efficiency
I Introduction
@@
The rise of Graph Neural Networks (GNNs) has brought about a transformative shift in the field of machine learning, particularly for tasks involving graph-structured data [1, 2, 3, 4]. These networks have demonstrated success across a wide array of real-world scenarios, including recommendation systems [5], traffic forecasting [6], chip design [7], and social network analytics[8]. The unique advantage of GNNs lies in their ability to effectively leverage the graph that represents the relationships between data points to optimize their feature representations. This enables GNNs to fully utilize the crucial graph topology inherent in graph-structured data, unlike traditional neural networks, which optimize data representations based solely on feature vectors, thus overlooking the valuable graph structure.
@@
However, like other neural networks, GNNs are susceptible to noise in the feature space[9, 10]. Additionally, the introduction of graph structure introduces another layer of interference, as GNNs are also affected by noise within the graph itself. Research has shown that even slight modifications to the graph—such as adding, removing, or rearranging edges—can significantly impact the performance of GNNs [11, 12]. The feature vectors contaminated by noise, along with erroneous and redundant edges in the graph topology, can be viewed as adversarial attacks that mislead the GNNs. To address this issue, researchers have proposed various methods to reduce the impact of noise on the model’s performance. These approaches can be divided into two categories: reactive and proactive defenses. Reactive defenses focus on identifying adversarial examples within the model’s inputs, as explored in previous studies[13, 14, 15]. On the other hand, proactive defenses aim to strengthen the model’s robustness, making it less susceptible to adversarial attacks, as demonstrated by methods proposed in works such as [16, 17].
@@
However, due to the diversity of noise types and model structures, it is challenging to develop effective methods for defending against noise across different scenarios. Unlike existing approaches that focus on mitigating the effects of noise, we argue that, given the inevitability of noise, the focus should be on how to better leverage it. In this paper, we propose a method that first identifies the training samples most vulnerable to noise from the dataset, and then uses these samples to construct a training set for training the GNN model. This approach effectively enhances the model’s ability to process noisy data correctly. Our experiments on multiple benchmark datasets demonstrate the effectiveness of this method.
II Preliminaries
II-A The Vulnerability of GNNs to Noise
@@
The vulnerability of Graph Neural Networks (GNNs) to noise refers to the difficulty GNNs face in maintaining consistent outputs when exposed to perturbations in the graph structure, such as the presence of incorrect or redundant edges in the input graph, or errors in the feature representations within the input feature matrix. [20] investigated the task- and model-agnostic vulnerabilities of GNNs, showing that these networks are very vulnerable to adversarial perturbations, regardless of the downstream task.
II-B Spectral Graph Theory
@@
According to spectral graph theory, when analyzing a graph, often only one or a few special eigenvalues and their corresponding eigenvectors are useful. For example, when analyzing the clustering structure of a graph using spectral graph theory, only the smallest eigenvalues and their corresponding eigenvectors are useful[18]. Therefore, for a Laplacian matrix, it is unnecessary to compute all the eigenpairs through direct eigen decomposition, and for a large-scale graph, the computational cost would be too high. The Courant-Fischer Minimax Theorem[19] allows the eigenvalues to be computed iteratively without explicitly calculating all the eigenvalues. Specifically:
@@
III Method
@@
Figure 1 shows the proposed method for training GNNs using non-robust training samples, which consists of four key phases: Phase (1) is GNN embedding, which uses the original input to train a GNN and takes the output of the hidden layer as the embedding result of the data’s feature matrix; In Phase (2), we use k-NN graphs to capture the underlying manifolds of the input feature matrix and the embedded feature matrix; In Phase (3), we calculate each node’s level of non-robustness using the input manifold and the output manifold; In Phase (4), we use the top non-robust samples selected from the training set to train the GNN. The details of these four phases are elaborated in the following sections.
III-A Phase 1: GNN Embedding
@@
To obtain the output data of the GNN model, we first train the GNN using the original feature matrix and input graph. Once trained, the GNN is then used to transform the input feature matrix. Since the final layer of a GNN is typically task-specific—such as a softmax layer for multi-class classification tasks—the outputs of the hidden layers prior to the final layer are used as embeddings of the original feature
III-B Phase 2: Manifold Capturing
The manifold of the data is typically captured using graphs, and k-nearest neighbor graphs are the most widely used for this purpose. In this phase, we use k-NN graphs to capture the underlying manifolds of the input feature matrix and the embedded feature matrix. The conventional kNN graph construction algorithm exhibits a time complexity of $O(|N|^{2})$, where $|N|$ is the total number of nodes in the graph. Such quadratic complexity makes it impractical for large-scale datasets. To overcome this limitation, an approximate kNN graph construction algorithm can be employed. This approach leverages an extension of the probabilistic skip list structure to achieve a reduced time complexity of $O(|N|\log|N|)$, which is more suitable for large-scale data processing [22].
III-C Phase 3: Adversarial Robustness Evaluation
Building on the theoretical framework introduced in [21], we leverage the largest generalized eigenvalues and their associated eigenvectors of $L_{Y}^{+}L_{X}$ to assess the stability of individual nodes in GNNs. Specifically, we calculate the weighted eigensubspace matrix $V_{s}\in\mathbb{R}^{|V|\times s}$ to perform spectral embedding of the input manifold $G_{X}=(V,E_{X})$, where $|V|$ represents the total number of nodes. The matrix $V_{s}$ is expressed as:
@@ III-C Phase 3: Adversarial Robustness Evaluation
Using $V_{s}$, we project the nodes of $G_{X}$ into an $s$-dimensional space, representing each node $p$ as the $p$-th row of $V_{s}$. The stability of an edge $(p,q)\in E_{X}$ is then quantified by computing the spectral embedding distance between nodes $p$ and $q$:
@@
A metric, referred to as the Spade score, is defined as the average embedding distance to the neighbors within the input manifold to assess the robustness of a node $p$:
@@
where $N_{X}(p)$ represents the set of neighbors of node $p$ in $G_{X}$. A higher $\text{Spade}(i)$ value suggests that node $i$ is more prone to adversarial attacks.
III-D Phase 4: GNN Training
@@
For the samples in the training set, we rank them in descending order based on their Spade scores and select the top portion to form a new training set. The points in this new training set are those that are highly susceptible to noise and likely to mislead the GNN. By training the GNN using only these noise-sensitive samples, we are effectively enhancing the GNN’s ability to resist being misled by such samples.
IV Experiments
IV-A Experimental Setup
@@
We use three standard citation network benchmark datasets to evaluate the proposed method: Cora, Citeseer, and Pubmed. In all these citation networks, nodes represent documents and edges correspond to citations. Each node is associated with a sparse bag-of-words feature vector and a class label.
@@
The Cora dataset consists of 2,708 scientific publications categorized into seven classes, forming a citation network with 5,429 links.
@@
The Citeseer dataset includes 3,327 scientific publications categorized into six classes, with a total of 4,732 links in the citation network. Compared to Cora, Citeseer is sparser and exhibits a more uneven class distribution, making it more challenging for graph-based classification methods.
The Pubmed dataset contains 19,717 scientific publications from the field of biomedicine, categorized into three classes. Each publication is represented by a Term Frequency-Inverse Document Frequency (TF-IDF) weighted word vector derived from a vocabulary of 500 words. The citation network in Pubmed includes 44,338 links, making it significantly larger than both Cora and Citeseer.
We extract 20% of the samples from the standard Cora dataset as the test set. From the remaining 80% of the samples, we construct two training sets: one by randomly selecting 80% of the samples and the other by selecting the top 80% non-robust nodes using our method. Both training sets are used to train GCNs with the same architecture and training parameters. To further ensure a fair comparison, we initialize both GCNs with the same random seed by setting torch.manual_seed(42), which guarantees identical random initialization of their parameters.
IV-B Experimental Results
IV-B1 Experimental Results on Graph Convolutional Networks (GCNs)
For the Cora dataset, the classification accuracy results on the test set as training epochs progress is shown in Figure 2 and Table I .

### v9

# Improving Graph Neural Network Training, Defense and Hypergraph Partitioning via Adversarial Robustness Evaluation

@@ Abstract
Graph Neural Networks (GNNs) are a highly effective neural network architecture for processing graph-structured data. Unlike traditional neural networks that rely solely on the features of the data as input, GNNs leverage both the graph structure, which represents the relationships between data points, and the feature matrix of the data to optimize their feature representation. This unique capability enables GNNs to achieve superior performance across various tasks. However, it also makes GNNs more susceptible to noise and adversarial attacks from both the graph structure and data features, which can significantly increase the training difficulty and degrade their performance. Similarly, a hypergraph is a highly complex structure, and partitioning a hypergraph is a challenging task. This paper leverages spectral adversarial robustness evaluation to effectively address key challenges in complex-graph algorithms. By using spectral adversarial robustness evaluation to distinguish robust nodes from non-robust ones and treating them differently, we propose a training-set construction strategy that improves the training quality of GNNs. In addition, we develop algorithms to enhance both the adversarial robustness of GNNs and the performance of hypergraph partitioning. Experimental results show that this series of methods is highly effective.
footnotetext: Correspondence to: yongyuw@mtu.edu
1 Introduction
@@
The rise of Graph Neural Networks (GNNs) has brought about a transformative shift in the field of machine learning, particularly for tasks involving graph-structured data [1, 2, 3, 4]. These networks have demonstrated success across a wide array of real-world scenarios, including recommendation systems [5], traffic forecasting [6], chip design [7], language processing [8] and social network analytics[9]. The unique advantage of GNNs lies in their ability to effectively leverage the graph that represents the relationships between data points to optimize their feature representations. This enables GNNs to fully utilize the crucial graph topology inherent in graph-structured data, unlike traditional neural networks, which optimize data representations based solely on feature vectors, thus overlooking the valuable graph structure.
However, like other neural networks, GNNs are susceptible to noise and adversarial attacks in the graph and feature space[10, 11]. . Research has shown that even slight modifications to the graph—such as adding, removing, or rearranging a small number of edges—can significantly impact the performance of GNNs [12, 13].
@@
To address this issue, researchers have proposed various methods to reduce the impact of noise on the model’s performance. These approaches can be divided into two categories: reactive and proactive defenses. Reactive defenses focus on identifying adversarial examples within the model’s inputs, as explored in previous studies[14, 15, 16]. On the other hand, proactive defenses aim to strengthen the model’s robustness, making it less susceptible to adversarial attacks, as demonstrated by methods proposed in works such as [17, 18].
@@
However, due to the diversity of noise types and model structures, developing effective methods to defend against noise across various scenarios remains a significant challenge. Unlike existing approaches that focus on mitigating the effects of noise, we argue that, given the inevitability of noise, the focus should shift toward how to better leverage it. From this new perspective, this paper proposes a framework that first identifies the most noise-sensitive training samples and then uses these samples to construct a smaller but more effective training set for training the GNN model. This approach enhances the model’s ability to correctly process noisy data.
In this work, we also propose a new approach to improve the defense ability of GNNs. Specifically, we apply spectral adversarial robustness evaluation to partition nodes in the graph into two groups: robust nodes that are relatively insensitive to noise and perturbations, and non-robust nodes that are more easily affected. We then treat these two groups differently. For robust nodes, we induce a subgraph containing only them from the original graph, and use a GNN to predict their class labels. For non-robust nodes, we leverage the topology of the original graph to propagate class information from robust nodes to obtain their labels
We also provide a efficient high performance hypergraph partitioning method based on spectral adversarial robustness evaluation and demonstrate its effectiveness on benchmark VLSI data sets.
2 Preliminaries
2.1 The Vulnerability of GNNs to noise and adversarial attacks
@@
The vulnerability of Graph Neural Networks (GNNs) to noise and adversarial attacks refers to the difficulty GNNs face in maintaining consistent outputs when exposed to perturbations in the graph structure, such as the presence of incorrect or redundant edges in the input graph, or errors in the feature representations within the input feature matrix. [23] investigated the task- and model-agnostic vulnerabilities of GNNs, showing that these networks are very vulnerable to adversarial perturbations, regardless of the downstream task.
2.2 Spectral Graph Theory
@@
According to spectral graph theory, analyzing a graph often requires only one or a few specific eigenvalues and their corresponding eigenvectors [20]. For example, when analyzing the clustering structure of a graph using spectral graph theory, only the smallest eigenvalues and their corresponding eigenvectors are useful[21]. Therefore, for a Laplacian matrix, it is unnecessary to compute all the eigenpairs through direct eigen decomposition, and for a large-scale graph, the computational cost would be too high. The Courant-Fischer Minimax Theorem[22] allows the eigenvalues to be computed iteratively without explicitly calculating all the eigenvalues. Specifically:
@@
3 Improving Graph Neural Network Training
3.1 Method
@@
Figure 1 shows the proposed method for training GNNs using top non-robust training samples in the training set, which consists of four key phases: Phase (1) is GNN embedding, which uses the original input to train a GNN and takes the output of the hidden layer as the embedding result of the data’s feature matrix; In Phase (2), we use k-NN graphs to capture the underlying manifolds of the input feature matrix and the embedded feature matrix; In Phase (3), we calculate each node’s level of non-robustness using the input manifold and the output manifold; In Phase (4), we use the top non-robust samples selected from the training set to train the GNN. In the rest of this section, we provide a detailed description of these four phases.
3.1.1 Phase 1: GNN Embedding
@@
To obtain the output data of the GNN model, we first train the GNN using the original feature matrix and input graph. Once trained, the GNN is then used to transform the input feature matrix. Since the final layer of a GNN is typically task-specific—such as a softmax layer for multi-class classification tasks—we take the output of the hidden layer before the final output layer as the embedding result.
3.1.2 Phase 2: Manifold Capturing
The manifold of the data is typically captured using graphs, and k-nearest neighbor graphs are the most widely used for this purpose. In this phase, we use k-nearest neighbor (k-NN) graphs to capture the underlying manifolds of the input feature matrix and the embedded feature matrix. The conventional k-NN graph construction algorithm exhibits a time complexity of $O(|N|^{2})$, where $|N|$ is the total number of nodes in the graph. Such quadratic complexity makes it impractical for large-scale datasets. To overcome this limitation, an approximate kNN graph construction algorithm can be employed [25]. This approach leverages an extension of the probabilistic skip list structure to achieve a reduced time complexity of $O(|N|\log|N|)$, which is more suitable for large-scale data processing.
3.1.3 Phase 3: Adversarial Robustness Evaluation
Let $L_{X}$ and $L_{Y}$ denote the Laplacian matrices corresponding to the k-NN graphs of the input feature matrix and the embedded feature matrix constructed in phase 2, respectively. Building on the theoretical framework introduced in [24], we leverage the largest generalized eigenvalues and their associated eigenvectors of $L_{Y}^{+}L_{X}$ to assess the robustness of each individual sample to noise during the processing by a given model. Specifically, we calculate the weighted eigensubspace matrix $V_{s}\in\mathbb{R}^{|V|\times s}$ to perform spectral embedding of the input manifold $G_{X}=(V,E_{X})$, where $|V|$ represents the total number of nodes. The matrix $V_{s}$ is expressed as:
@@ 3.1.3 Phase 3: Adversarial Robustness Evaluation
Using $V_{s}$, we project the nodes of $G_{X}$ into an $s$-dimensional space, representing each node $p$ as the $p$-th row of $V_{s}$. According to [24], the stability of an edge $(p,q)\in E_{X}$ is quantified by computing the spectral embedding distance between nodes $p$ and $q$:
@@
To assess the robustness of a node $p$, a metric referred to as the Spade score is defined as the average embedding distance to the neighbors within the input manifold:
@@
where $N_{X}(p)$ represents the set of neighbors of node $p$ in $G_{X}$. A higher $\text{Spade}(p)$ value suggests that node $p$ is more prone to adversarial attacks.
3.1.4 Phase 4: GNN Training
@@
For the samples in the training set, we rank them in descending order based on their Spade scores and select the top portion to form a new, smaller training set. The points in this new training set are those that are highly susceptible to noise and likely to mislead the GNN. By training the GNN using only these noise-sensitive samples, we are effectively enhancing the GNN’s ability to resist being misled by such samples.
3.2 Experiments
We have evaluated our framework on three of the most popular GNN models and three classic benchmark datasets.
3.2.1 Data Sets
@@
We evaluate the proposed method on three standard citation network benchmarks, where nodes represent documents, edges correspond to citations, and each node is associated with a sparse bag-of-words feature vector and a class label. The key statistics of these datasets are summarized in Table 1.
@@
  • Cora: A dataset of scientific publications categorized into seven classes.
@@
  • Citeseer: A dataset of scientific publications that is sparser than Cora and exhibits a more uneven class distribution, making it more challenging for graph-based classification methods.
  • PubMed: A large biomedical dataset where nodes are represented by TF-IDF word vectors from a 500-word vocabulary.
Table 1: Statistics of datasets used in our experiments.
Dataset Nodes Edges Classes Features
Cora 2,708 5,429 7 1,433
Citeseer 3,327 4,732 6 3,703
PubMed 19,717 44,338 3 500
3.2.2 GNN Models
We use the following three commonly used GNN models for our experiments:
  • Graph Convolutional Network (GCN): A widely used model that applies convolution-like operations to graph-structured data by aggregating information from neighboring nodes using a layer-wise propagation rule.
  • Graph Attention Network (GAT): Introduces an attention mechanism to assign different importance weights to neighboring nodes during the message-passing process.
  • GraphSAGE: An inductive model that learns aggregation functions to generate node embeddings by sampling and aggregating features from a fixed number of neighboring nodes.
3.2.3 Detailed Model Implementation and Parameter Settings
  • GCN: We use the same GCN architecture and parameter settings for all three datasets. The model consists of two graph convolutional layers: the first layer transforms node features to a 64-dimensional space, and the second layer maps them to the number of output classes, with ReLU activation applied after the first convolution. For the Cora dataset, the learning rate is set to 0.01. For the Citeseer and PubMed datasets, it is set to 0.1.
  • GAT: For the Cora dataset, the GAT model consists of two layers: the first layer has 4 attention heads with 8 hidden units each, followed by an ELU activation, and the second layer has 1 attention head that outputs the number of class logits. For the Citeseer dataset, the GAT model also has two layers: the first layer uses 8 attention heads with 8 hidden units each, followed by an ELU activation, and the second layer has 1 attention head that outputs the class logits. For the PubMed dataset, the GAT model has the same structure as the Citeseer model but includes layer normalization after each layer and uses a Leaky ReLU activation after the first layer. For the Cora dataset, the learning rate is set to 0.01. For the Citeseer dataset, it is set to 0.1, and for the PubMed dataset, it is set to 0.05.
  • GraphSAGE: The GraphSAGE architecture for the Cora and Citeseer datasets consists of two layers: the first layer uses SAGEConv with 64 hidden units, followed by batch normalization, ReLU activation, and a dropout layer with a probability of 0.5. The second layer uses SAGEConv with 64 input features and outputs the class labels. The architecture for the PubMed dataset also consists of two layers, but the first layer uses SAGEConv with 32 hidden units, followed by batch normalization, ReLU activation, and a dropout layer. The second layer uses SAGEConv with 32 input features and outputs the class labels. For all three datasets, the learning rate is set to 0.05.
For all the experiments, the Adam optimizer is used with a weight decay of $5\times 10^{-4}$, and the loss function for multi-class classification is CrossEntropyLoss. The number of eigenvectors is set to 10, and the value of $k$ in k-NN is set to 20.
3.2.4 Results
We first perform an 80-20 split of the entire dataset by randomly sampling 80% of the samples to form the original full training set and using the remaining 20% as the test set. Then, we construct two different types of training sets: (1) forming the training set by selecting the top 80% of non-robust samples from the original full training set using our method; (2) forming the training set by randomly sampling 80% of the original full training set.
We train the GNN models using the original full training set and the two smaller training sets mentioned above, and the classification accuracy across different epochs is shown in Table 2. In these charts, Random represents the training set constructed by randomly sampling 80% of the original full training set, Ours represents the training set constructed using our method, and Full represents the original full training set.
From the experimental results, it can be observed that in all 9 combinations of the three GNN models and the three datasets, the training set constructed using our method achieved the highest classification accuracy after 5 epochs of training. Moreover, the advantage of our method is particularly evident on the largest dataset, PubMed. On PubMed, our method outperforms the second-best method by 14.1%, 7.4%, and 3.4% in accuracy for the GCN, GAT, and GraphSAGE models, respectively.
Our method not only significantly outperforms the randomly sampled training sets of the same size but also demonstrates substantial advantages compared to the full training sets, which are 20% larger, highlighting the effectiveness of our approach. The experimental results further show that randomly selecting 80% of the training samples from the full training set yields almost no difference in performance compared to using the full set. However, when 80% of the samples are selected using our method, which prioritizes the most non-robust nodes, the training performance is significantly improved. This demonstrates that focusing on non-robust samples during training, rather than uniformly sampling all nodes, enables the model to better adapt to unavoidable noise, resulting in superior performance on the test set.